OpenStudy (anonymous):

x=2y+7 x=y+4 solve please

OpenStudy (anonymous):

x = 1 y = 3

jimthompson5910 (jim_thompson5910):

x=y+4 2y+7=y+4 2y=y+4-7 2y-y=4-7 y=4-7 y=-3 -------------------------- x = y+4 x = -3+4 x = 1 So the solutions are x = 1 and y = -3

OpenStudy (anonymous):

@jim_thompson5910, you are very helpful! Thanks!

jimthompson5910 (jim_thompson5910):

yw, glad to be of help